one leg of a triangle is 14ft longer than the other leg. The legnth of the hypotenuse is 26ft. Find the lengths of the other 2 sides.
Answer by richwmiller(17219)

a^2+b^2=26^2
a=b+14
(b+14)^2+b^2=26^2
solve for b
accept only positive answers.